OpenStudy (anonymous):

Find the non-extraneous solutions of the square root of the quantity x plus 7 minus 4 equals quantity x plus 3

OpenStudy (anonymous):

(x + 3)(1/2) = x + 3 x + 3 = x^2 + 9 x^2 -x - 6 = 0 (x - 3) (x+2) = 0 x is +3 or -2 no idea what an extraneous solution is

OpenStudy (anonymous):

extraneous is a solution of a simplified version of an equation that does not satisfy the original equation.